Determine the order of the reactant. aA bB Time [A] (s) 0 0.0100 4 0.0071 0.0055 8 12 0.0045 16 0.0038
Answer #1

Firstly we have to check whether the reaction is first order or not by using relation k1 = 1/t ln a/(a-x)

Calculated k1 is not constant in all cases.. Therefore this reaction will not be first order.

Time [A] 0.01 a 6.0071 7 0.0055 0.0045 a-x 0.0038 14 118 112 16 - 2= a-la-x) 0:0) - 0.0071 -0.0029 0.01 -0.0055 0.0045 0.01 -